I am struggling to understand how can 3/4-2/3=1/12
But -3/4-2/4= 17/12
I know it is prob a simple answer but I am at a loss.

Answer by ptaylor(2198)

ok
3/4 - 2/3= The lowest common denominator is 12
3/4=9/12
and
2/3=8/12
soooo
9/12-8/12=1/12
-3/4-2/4=
-9/12-8/12=-17/12
Hope this helps---ptaylor